A student walks across campus from the residence hall to the science lab in 18 minutes. The science lab is 0.3 km from the residence hall. What is the student's average speed in km/h? (Note: Give the answer without a unit)


1
Expert's answer
2021-02-03T16:15:23-0500

First, find out what fraction of an hour 18 minutes are: 18/60=0.3 h.

The average speed in km/h is


"s=\frac{d}{t}=\frac{0.3}{0.3}=1\text{ km/h}."